OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 5x = x^2 - 24

OpenStudy (anonymous):

First bring 5x over to the right. y = x^2 - 5x - 24 Factor y = (x+3)(x-8) Set y = 0 (x+3)(x-8) = 0 x = -3 or x = 8 The x-intercepts are therefore (-3, 0) and (0, 8). :)

OpenStudy (saifoo.khan):

Goddess of mars to Rescue.

OpenStudy (anonymous):

xD It is macra, plural of macron, which is a overbar in some languages to indicate a stressed vowel. As is, āēīōū. :)

OpenStudy (anonymous):

but mars is great too lol

OpenStudy (anonymous):

OK this girl deservess like some sort of level up to like 30! she is just so fast!

OpenStudy (anonymous):

X) Thank you, I am very fast at typing

OpenStudy (saifoo.khan):

MACRA = http://www.macra.org.my/welcome.php

OpenStudy (anonymous):

Your answers arew quick, and correct. I hope your on all day xD

OpenStudy (anonymous):

hahahahahaha

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

haha :) I will be on for awhile. But doing language homework too, so in and out lol

OpenStudy (saifoo.khan):

so you are Goddess of Malaysian Air-Conditioning & Refrigeration Association.

OpenStudy (anonymous):

XD I should really write that down haha

OpenStudy (saifoo.khan):

ROFL.

OpenStudy (anonymous):

Ok, cause im doing a semester exam, and its fourty two questions

OpenStudy (anonymous):

What?

OpenStudy (saifoo.khan):

ahh, no leave it.

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

haha lag in messages nevermind

OpenStudy (saifoo.khan):

this website sucks at lagging.

OpenStudy (anonymous):

Yeah, and the times that people posted. Like it says your last message is from 2 minutes ago lol XD

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

okay i'll be back in a bit, as now i have to attend to my own hw lol

OpenStudy (saifoo.khan):

haha, okay!!